1/1

Ernst Anschütz,

Fox, you stole my goosey gander

tradukita de N. N. 24

Fox, you stole my goosey gander, |: Give him back to me, :| |: Else I’ll call the forest ranger Who’ll shoot you with glee. :|

He has a long and big shot gun |: He’ll aim at your head, :| |: Down your face, red ink will run And then you will be dead. :|

Dearest fox, let me advise you |: A thief not to be, :| |: With a mouse you may make do, But leave the goose for me. :|

Traduko de la Germana poemo “Fuchs, du hast die Gans gestohlen” de Ernst Anschütz (∗1780-10-28 – †1861-12-18) en la Anglan de N. N. 24.

Arg-563-1137 (2010-09-19 17:27:27)

Mi, Manfred Retzlaff, trovis tiun ĉi angligon de la kanto en http: // ingeb. org/ Lieder/ FuchsDuH. html .

http://www.poezio.net/version?poem-id=563&version-id=1137